OpenStudy (anonymous):

solve this system as an ordered pair 2x-y=-1 3x+5y=18

OpenStudy (anonymous):

\(\Large 2x-y=-1\Rightarrow y=2x+1\) instead of the value y into rest equation Therefore, \(\Large 3x+5(2x+1)=18\\\Large \Leftrightarrow 13x=13 \Rightarrow x=1\\\Large \Rightarrow y=2x+1=2+1=3\)